# Second Law of Thermodynamics without Einstein Relation

### v2

@@ Abstract
Materials that are constantly driven out of thermodynamic equilibrium, such as active and living systems, typically violate the Einstein relation. This may arise from active contributions to particle fluctuations which are unrelated to the dissipative resistance of the surrounding medium. We show that in these cases the widely used relation between informatic entropy production and heat dissipation does not hold. Consequently, fluctuation relations for the mechanical work, such as the Jarzynski and Crooks theorems, are invalid. We relate the breaking of the correspondence between entropy production and heat dissipation to departure from the fluctuation-dissipation theorem. We propose a temperature-like variable which restores this correspondence and gives rise to a generalized second law of thermodynamics, whereby the dissipated heat is necessarily nonnegative and vanishes at equilibrium. The Clausius inequality, Carnot maximum efficiency theorem, and relation between the extractable work and the change of free energy are recovered as well.
Classical thermodynamics is the central theory for the possible forms of energy transfer in materials. Its many useful consequences, including the Carnot efficiency theorem and the bounds on extractable work via changes in free energy, establish nontrivial and universal relations based on a few macroscopic observables. Among these variables is the temperature, which ties microscopic statistics of equilibrium systems to their energetics, e.g., through the equipartition principle or by imposing the mean energy as a constraint for entropy maximization. In light of its vast applicability for characterizing the behaviors of equilibrium materials, there is an ongoing effort to extend thermodynamics to systems that are constantly driven out of equilibrium, for example, self-propelled particles [1, 2, 3], sustained chemical reactions [4, 5], and living systems [6]. Among the complications associated with nonequilibrium systems is the inadequacy of temperature and other variables as thermodynamic state functions [7, 8, 9, 10, 11, 12].
One of the consequences of thermodynamic equilibrium, where the system’s configurations are Boltzmann-distributed, is the Einstein relation (ER) [13, 14],
@@
The diffusivity $\mathbf{D}$ and mobility $\boldsymbol{\mu}$ characterize, respectively, the strength of thermal fluctuations and dissipative attenuation of motion. In general they may be space-dependent matrices, reflecting translational and rotational symmetry breaking. At equilibrium the effects of fluctuations and dissipation exactly balance [14], leading to Eq. (1). However, in cases where forces are not conservative [8, 15], or if fluctuations arise from coarse-graining of fast athermal or active forces [16, 17, 18, 19, 20, 10, 21, 22, 23, 24], there should not be a generic relation between fluctuations (diffusivity) and dissipation (mobility). Thus, more often than not, in far-from-equilibrium cases such as active biopolymer networks [25, 26, 27], bacterial swarms [28, 29], and driven chemical reactions [30, 5], the ER does not hold. In cases where $\mathbf{D}$ is proportional to $\boldsymbol{\mu}$, one can use Eq. (1) to define an effective temperature [7, 10, 31, 19, 18]. However, in the general case, these two position-dependent tensors are not related by a constant uniform scalar, and Eq. (1) is violated.
@@
One of the theories addressing nonequilibrium systems is stochastic thermodynamics [32]. By considering Langevin equations or discrete Markov chains it establishes trajectory-wise energy conservation (leading to the first law of thermodynamics) [33, 34, 35] and gives a wealth of fluctuation relations (FRs) [36], such as the Jarzynski and Crooks relations [37, 38, 39]. The latter also yields the second law of thermodynamics for the average entropy changes and heat increments. Yet, implicitly, all the aforementioned FRs and extended laws of thermodynamics have relied on the existence of a temperature, defined through the ER, to connect microscopic statistics to thermodynamic energy, mechanical and chemical work, and calorimetric heat.
@@
Several definitions of a nonequilibrium temperature were proposed based on generalizations of the fluctuation-dissipation theorem (FDT) and FRs in active matter [10, 20, 40, 41], systems with memory [16, 17], and anomalous diffusion [42, 43]. Most of these require a thermal bath that obeys the ER. In this Letter, we adopt an alternative approach, which is not derived from extentions of FRs or FDT, to identify a temperature-like variable directly, such that the classical-thermodynamic results are recovered.
@@
In some cases, the ER is required to keep the extended FRs connected to physical measurables. For example, the Hatano-Sasa relation [44], an extension of the Jarzynski equality, requires Markovian dynamics only. It therefore relaxes Jarzynski’s assumption of a steady-state Boltzmann distribution for a given value of the external protocol [38], thus relaxing also the ER requirement. Yet, the effective Hamiltonian and work used by this relation coincide with their actual, measurable counterparts only under those extra assumptions. Otherwise, they demand knowledge of the complete microstates’ steady-state distribution, which is usually inaccessible or requires simplistic models [45]. The temperature-like variable proposed here, arising directly from thermodynamics, does not rely on those assumptions.
@@
Overdamped stochastic energetics.—We consider an ensemble of $N$ interacting degrees of freedom (particles), subjected to a deterministic external time-dependent protocol, $\Lambda_{t}$. The time-dependent stochastic microstate is an $N$-component vector, $\mathbf{X}_{t}$. We couple the particles to an arbitrary medium determining the mobility $\boldsymbol{\mu}(\mathbf{x},\lambda)$ and diffusivity $\mathbf{D}(\mathbf{x},\lambda)$, such that the particles follow the overdamped Langevin equation [Eq. (2) below]; no other property of the surrounding medium is specified. The aforementioned $\mathbf{X}_{t}$ and $\Lambda_{t}$ denote a particular time-dependent realizations of $\mathbf{x}$ and $\lambda$. The particles are subjected to an additional external force $\mathbf{F}_{\mathrm{ext}}(\mathbf{x},\lambda)$, which may be nonconservative and is excluded from the internal energy. We denote the total ($N$-component) force acting on the particles by $\mathbf{F}=\mathbf{F}_{\mathrm{ext}}-\boldsymbol{\nabla}H$, where $H(\mathbf{x},\lambda)$ is the system’s Hamiltonian. The system is driven out of equilibrium due to the nonconservative force $\mathbf{F}_{\mathrm{ext}}$ and/or violation of ER [Eq. (1)]. Thus, the system’s steady-state distribution is not Boltzmann, nor can a temperature be conventionally identified.
@@
where $\mathbf{b}$ is the noise magnitute satisfying $\mathbf{D}=\mathbf{b}\cdot\mathbf{b}^{\mathrm{T}}/2$, $\boldsymbol{\eta}_{t}$ is the normalized thermal noise [$\langle\boldsymbol{\eta}_{t}\rangle=\mathbf{0}$, $\langle\boldsymbol{\eta}_{t}\boldsymbol{\eta}_{t^{\prime}}\rangle=\mathbf{I}\delta(t-t^{\prime})$ where $\mathbf{I}$ is the identity matrix], and the equation is written in the Itô convention [46]. The “spurious drift” term ${\boldsymbol{\nabla}}\cdot\mathbf{D}$ ensures that Boltzmann distribution is obtained at equilibrium, i.e., in systems with a well-defined temperature, Eq. (1), under conservative forces [47, 48]. The corresponding Fokker-Planck equation [49] for the microstates distribution $p(\mathbf{x},t)$ is
@@
We follow Sekimoto’s approach [33, 34] to identify the internal energy, work, and heat for trajectories in generic Langevin systems. The time-derivative of the Hamiltonian is [46]
@@
The above is a natural adaptation of the internal energy, work, and heat to the microscale [33, 34]. A similar extension of the system’s entropy is less trivial. Consider the instantaneous nonequilibrium stochastic entropy [32, 36], $S(\mathbf{x},t)=-\ln p(\mathbf{x},t)$, whose ensemble average is the Shannon entropy [51, 52, 53], $\langle S\rangle=-\int{\text{d}}\mathbf{x}\,p(\mathbf{x},t)\ln p(\mathbf{x},t)$. The relation between classical-thermodynamic entropy and stochastic entropy is established for equilibrium states by finding the distribution $p$ that maximizes $\langle S\rangle$ [54] (for a fixed protocol value $\lambda$). Combining the stochastic chain rule [46] and Eqs. (2)–(4) yields the time-derivative of $S_{t}=S(\mathbf{X}_{t},t)$,
@@
If a well-defined temperature $T$ exists, it is instructive to consider the thermodynamic heat dissipation, $k_{\mathrm{B}}TR_{t}$. It quantifies how much extra calorimetric heat $\dot{Q}_{t}$ is wasted in an irreversible process. According to the second law of thermodynamics, the average thermodynamic dissipation $\langle\dot{R}_{t}\rangle$ is nonnegative [56],
@@
Informatic entropy production (IEP).—The IEP, $\Sigma_{t}$, quantifies the breaking of time-reversal symmetry in the microstates’ dynamics [32]. Consider a discretized trajectory drawn from Eq. (2), $(\mathbf{X}_{0},\dots,\mathbf{X}_{K\Delta t})=(\mathbf{x}_{0},\dots,\mathbf{x}_{K})$, with arbitrarily-small timestep $\Delta t$. The probability distribution of obtaining this trajectory under the protocol sequence $(\lambda_{0},\lambda_{1},\ldots,\lambda_{K})$, $\overrightarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{K})$, is compared with the probability distribution of obtaining the reverted trajectory, $\overleftarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{K},\mathbf{x}_{K-1},\ldots,\mathbf{x}_{0})$, under the reverted protocol $(\lambda_{K},\lambda_{K-1},\ldots,\lambda_{0})$. Explicitly [36, 32],
@@
In Eq. (12) we identify the time-derivative of the entropy, $\dot{S}_{t}=\ln[p(\mathbf{x}_{k+1},(k+1)\Delta t)/p(\mathbf{x}_{k},k\Delta t)]/\Delta t$, given by Eq. (9). We refer to the remainder, $\Omega_{t}$, as the “informatic heat” (IH). The propagator in the denominator, $\Pr(\mathbf{x}_{k},\lambda_{k}|\mathbf{x}_{k+1},\lambda_{k+1})$, is the exact same propagator of Eq. (2), where the jump process is reverted [i.e., the process starts at $\mathbf{x}_{k+1}$ at time $t=k\Delta t$ and ends up at $\mathbf{x}_{k}$ at time $(k+1)\Delta t$ under the protocol value $\lambda_{k+1}$].
@@
Therefore, if $\mathbf{D}$ and $\boldsymbol{\mu}$ are not related by a scalar, the informatic $\dot{\Sigma}_{t}$ is no longer related to the measurable $\dot{R}_{t}$, even on average. Consequently, the nonnegativity of $\langle\dot{R}_{t}\rangle$ (i.e., the second law) can no longer be established based on trajectory statistics. Below we propose a definition of a nonequilibrium temperature that mends this discrepancy.
@@
We now impose the following physical conditions on Eq. (14): (i) To satisfy the second law, $\langle\dot{R}_{t}\rangle$ should be a nonnegative functional of $\mathbf{J}$. (ii) To agree with Onsager’s theory [61, 62, 63], the right-hand side should reduce to a quadratic form of $\mathbf{J}$ near equilibrium and turn the second-law inequality into an equality for reversible processes ($\mathbf{J}\to\mathbf{0}$). (iii) The variable $\mathcal{T}$ should be a configuration-independent scalar. The simplest choice for the dissipation functional that satisfies these three conditions is $\langle\dot{R}_{t}\rangle=\int{\text{d}}\mathbf{x}(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J})/p$ [64]. To complete our construction, we solve the equation $\langle\dot{R}_{t}\rangle=\langle\dot{S}_{t}\rangle-\langle\dot{Q}_{t}\rangle/(k_{\mathrm{B}}\mathcal{T})$ for $k_{\mathrm{B}}\mathcal{T}$, and find the following drive-dependent generalized temperature:
@@
We have thus recovered the second law of thermodynamics.
Equation (15) is our central result. If ER holds, such that $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}=(k_{\mathrm{B}}T)^{-1}$, we readily get $\mathcal{T}=T$. Otherwise, the generalized temperature is an intricate quantity. This is expected for nonequilibrium systems with arbitrarily complex $\mathbf{J}$, $\mathbf{D}$, $\boldsymbol{\mu}$, and $\mathbf{F}$. Furthermore, $\mathcal{T}$ depends on time via the protocol $\Lambda_{t}$ affecting all these quantities. Even for a constant protocol $\Lambda_{t}=\mathrm{const}$, $\mathcal{T}$ will change in time due to the instantaneous distribution $p(\mathbf{x},t)$ entering $\mathbf{J}$. Indeed, nonequilibrium steady states generally have $\mathbf{J}\neq\mathbf{0}$, as nonconservative forces and drift create persistent fluxes across scales. Despite these caveats, Eq. (15) should play an experimentally meaningful role as a generalized temperature in nonequilibrium thermodynamics [21, 31, 19, 18], as discussed below.
@@
which, combined with Eq. (9) and the relation $\dot{\Sigma}_{t}=\dot{S}_{t}+\dot{\Omega}_{t}$, gives the IEP rate as
@@
Note that the correspondence holds for the means only and not for the full statistics, $\dot{Q}_{t}\neq-k_{\mathrm{B}}\mathcal{T}\dot{\Omega}_{t}$ and $\dot{R}_{t}\neq\dot{\Sigma}_{t}$. For example, the variances of the former are different, $\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})p{\text{d}}\mathbf{x}$, whereas $\langle({\text{d}}Q_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int\mathbf{F}\cdot\mathbf{D}\cdot\mathbf{F}p{\text{d}}\mathbf{x}\neq(k_{\mathrm{B}}\mathcal{T})^{2}\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}$. (The subscript ${\mathrm{c}}$ denotes a cumulant.) Since the Jarzynski equality [37, 38] and the Crooks fluctuation theorem [39] require statistics beyond the mean, they are not restored by the generalized $\mathcal{T}$. Thus, without the ER, the various FRs [36] remain related only to informatic quantities (the IEP and IH) [44] and not to thermodynamics (dissipation and heat).
@@
Nevertheless, generalized versions of the FDT and the Harada-Sasa relation can be derived in the absence of the ER. Here we outline these results; see the Supplementary Material (SM) for details [67]. Instead of considering the response to an impulsive linear potential gradient as was done in Ref. [65], we consider a perturbation of the force of the form $\delta\mathbf{F}(\mathbf{x},s)=-\delta h_{\mu}(s)\boldsymbol{\mu}^{-1}(\mathbf{x},s)\cdot\mathbf{D}(\mathbf{x},s)\cdot{\boldsymbol{\nabla}}x_{\mu}$ at time $s<t$, in response to the external perturbation $\delta\mathbf{h}(s)$. The difference between the response of the average force to such perturbation and its correlation with the position are found to be related to the thermodynamic heat [67],
@@
$\langle\dot{\Omega}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.+\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}](\mathbf{X}_{t},t)\rangle\right)\,.$ (21)

### v3

@@ Abstract
Materials that are constantly driven out of thermodynamic equilibrium, such as active and living systems, typically violate the Einstein relation. This may arise from active contributions to particle fluctuations which are unrelated to the dissipative resistance of the surrounding medium. We show that in these cases the widely used relation between informatic entropy production and heat dissipation does not hold. Consequently, fluctuation relations for the mechanical work, such as the Jarzynski and Crooks theorems, are invalid. We relate the breaking of the correspondence between entropy production and heat dissipation to departure from the fluctuation-dissipation theorem. We propose a temperaturelike variable that restores this correspondence and gives rise to a generalized second law of thermodynamics, whereby the dissipated heat is necessarily non-negative and vanishes at equilibrium. The Clausius inequality, Carnot maximum efficiency theorem, and relation between the extractable work and the change of free energy are recovered as well.
Classical thermodynamics is the central theory for the possible forms of energy transfer in materials. Its many useful consequences, including the Carnot efficiency theorem and the bounds on extractable work via changes in free energy, establish nontrivial and universal relations based on a few macroscopic observables. Among these variables is the temperature, which ties microscopic statistics of equilibrium systems to their energetics, e.g., through the equipartition principle or by imposing the mean energy as a constraint for entropy maximization.
In light of its vast applicability for characterizing the behaviors of equilibrium materials, there is an ongoing effort to extend thermodynamics to systems that are constantly driven out of equilibrium, for example, self-propelled particles [1, 2, 3], sustained chemical reactions [4, 5], and living systems [6]. Among the complications associated with nonequilibrium systems is the inadequacy of temperature and other variables as thermodynamic state functions [7, 8, 9, 10, 11, 12].
One of the consequences of thermodynamic equilibrium, where the system’s configurations are Boltzmann distributed, is the Einstein relation (ER) [13, 14],
@@
The diffusivity $\mathbf{D}$ and mobility $\boldsymbol{\mu}$ characterize, respectively, the strength of thermal fluctuations and dissipative attenuation of motion. In general they may be space-dependent matrices, reflecting translational and rotational symmetry breaking. At equilibrium with a bath of temperature $T$, the effects of fluctuations and dissipation exactly balance [14], leading to Eq. (1). Systems away from equilibrium, such as supercooled liquids [15] and systems under external forces [8], typically do not satisfy Eq. (1) with the ambient temperature $T$. In such cases Eq. (1) has been used to define an effective temperature [7, 10, 16, 17, 18]. However, if the forces are not conservative [8, 19], or when fluctuations arise from coarse graining of fast athermal or active forces [20, 21, 18, 17, 22, 10, 23, 24, 25, 26], there is generally no relation between the two position-dependent tensors $\mathbf{D}$ and $\boldsymbol{\mu}$. Thus, in far-from-equilibrium cases such as active biopolymer networks [27, 28, 29], bacterial swarms [30, 31], and driven chemical reactions [32, 5], the ER is invalid, as no scalar $T$ exists to satisfy Eq. (1). In this Letter, we propose a temperaturelike variable for systems exhibiting this severe breakdown of the Einstein relation.
@@
One of the theories addressing nonequilibrium systems is stochastic thermodynamics [33]. By considering Langevin equations or discrete Markov chains, it establishes trajectorywise energy conservation (leading to the first law of thermodynamics) [34, 35, 36] and gives a wealth of fluctuation relations (FRs) [37], such as the Jarzynski and Crooks relations [38, 39, 40]. The latter also yields the second law of thermodynamics for the average entropy changes and heat increments. Yet, implicitly, all the aforementioned FRs and extended second law of thermodynamics have relied on the existence of a temperature, defined through the ER, to connect microscopic statistics to thermodynamic energy, mechanical and chemical work, and calorimetric heat.
@@
Several definitions of a nonequilibrium temperature were proposed based on generalizations of the fluctuation-dissipation theorem (FDT) and FRs in active matter [10, 22, 41, 42], systems with memory [20, 21], and anomalous diffusion [43, 44]. Most of these require a thermal bath that obeys the ER. In this Letter, we adopt an alternative approach, which is not derived from extensions of FRs or FDT, to identify a temperaturelike variable directly, such that the classical-thermodynamic results are recovered.
@@
In some cases, the ER is required to keep the extended FRs connected to physical measurables. For example, the Hatano-Sasa relation [45], an extension of the Jarzynski equality, requires Markovian dynamics only. It therefore relaxes Jarzynski’s assumption of a steady-state Boltzmann distribution for a given value of the external protocol [39], thus relaxing also the ER requirement. Yet, the effective Hamiltonian and work used by this relation coincide with their actual, measurable counterparts only under those extra assumptions. Otherwise, they demand knowledge of the complete microstates’ steady-state distribution, which is usually inaccessible or requires simplistic models [46]. The temperaturelike variable proposed here, arising directly from thermodynamics, does not rely on those assumptions.
@@
Overdamped stochastic energetics.—We consider an ensemble of $N$ interacting degrees of freedom (particles), subjected to a deterministic external time-dependent protocol, $\Lambda_{t}$. The time-dependent stochastic microstate is an $N$-component vector, $\mathbf{X}_{t}$. We couple the particles to an arbitrary medium determining the mobility $\boldsymbol{\mu}(\mathbf{x},\lambda)$ and diffusivity $\mathbf{D}(\mathbf{x},\lambda)$, such that the particles follow the overdamped Langevin equation [Eq. (2) below]; no other property of the surrounding medium is specified. The aforementioned $\mathbf{X}_{t}$ and $\Lambda_{t}$ denote a particular time-dependent realizations of $\mathbf{x}$ and $\lambda$. The particles are subjected to an additional external force $\mathbf{F}_{\mathrm{ext}}(\mathbf{x},\lambda)$, which may be nonconservative and is excluded from the internal energy. We denote the total ($N$-component) force acting on the particles by $\mathbf{F}=\mathbf{F}_{\mathrm{ext}}-\boldsymbol{\nabla}H$, where $H(\mathbf{x},\lambda)$ is the system’s Hamiltonian. The system is driven out of equilibrium due to the nonconservative force $\mathbf{F}_{\mathrm{ext}}$ [47] and/or violation of ER [Eq. (1)]. Thus, the system’s steady-state distribution is not Boltzmann, nor can a temperature be conventionally identified.
@@
where $\mathbf{b}$ is the noise magnitude satisfying $\mathbf{D}=\mathbf{b}\cdot\mathbf{b}^{\mathrm{T}}/2$, $\boldsymbol{\eta}_{t}$ is the normalized thermal noise [$\langle\boldsymbol{\eta}_{t}\rangle=\mathbf{0}$, $\langle\boldsymbol{\eta}_{t}\boldsymbol{\eta}_{t^{\prime}}\rangle=\mathbf{I}\delta(t-t^{\prime})$, where $\mathbf{I}$ is the identity matrix], and the equation is written in the Itô convention [48]. The “spurious drift” term ${\boldsymbol{\nabla}}\cdot\mathbf{D}$ ensures that Boltzmann distribution is obtained at equilibrium, i.e., in systems with a well-defined temperature, Eq. (1), under conservative forces [49, 50]. The corresponding Fokker-Planck equation [51] for the microstates distribution $p(\mathbf{x},t)$ is
@@
We follow Sekimoto’s approach [34, 35] to identify the internal energy, work, and heat for trajectories in generic Langevin systems. The time derivative of the Hamiltonian is [48]
@@
The above is a natural adaptation of the internal energy, work, and heat to the microscale [34, 35]. A similar extension of the system’s entropy is less trivial. Consider the instantaneous nonequilibrium stochastic entropy [33, 37], $S(\mathbf{x},t)=-\ln p(\mathbf{x},t)$, whose ensemble average is the Shannon entropy [53, 54, 55], $\langle S\rangle=-\int{\text{d}}\mathbf{x}\,p(\mathbf{x},t)\ln p(\mathbf{x},t)$. The relation between classical-thermodynamic entropy and stochastic entropy is established for equilibrium states by finding the distribution $p$ that maximizes $\langle S\rangle$ [56] (for a fixed protocol value $\lambda$). Combining the stochastic chain rule [48] and Eqs. (2)–(4) yields the time derivative of $S_{t}=S(\mathbf{X}_{t},t)$,
@@
If a well-defined temperature $T$ exists, it is instructive to consider the thermodynamic heat dissipation, $k_{\mathrm{B}}TR_{t}$. It quantifies how much extra calorimetric heat $\dot{Q}_{t}$ is wasted in an irreversible process. According to the second law of thermodynamics, the average thermodynamic dissipation $\langle\dot{R}_{t}\rangle$ is non-negative [58],
@@
Informatic entropy production (IEP).—The IEP, $\Sigma_{t}$, quantifies the breaking of time-reversal symmetry in the microstates’ dynamics [33]. Consider a discretized trajectory drawn from Eq. (2), $(\mathbf{X}_{0},\dots,\mathbf{X}_{K\Delta t})=(\mathbf{x}_{0},\dots,\mathbf{x}_{K})$, with arbitrarily small time step $\Delta t$. The probability distribution of obtaining this trajectory under the protocol sequence $(\lambda_{0},\lambda_{1},\ldots,\lambda_{K})$, $\overrightarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{0},\mathbf{x}_{1},\ldots,\mathbf{x}_{K})$, is compared with the probability distribution of obtaining the reverted trajectory, $\overleftarrow{\mathcal{P}}\equiv\Pr(\mathbf{x}_{K},\mathbf{x}_{K-1},\ldots,\mathbf{x}_{0})$, under the reverted protocol $(\lambda_{K},\lambda_{K-1},\ldots,\lambda_{0})$. Explicitly [37, 33],
@@
In Eq. (12) we identify the time derivative of the entropy, $\dot{S}_{t}=\ln[p(\mathbf{x}_{k+1},(k+1)\Delta t)/p(\mathbf{x}_{k},k\Delta t)]/\Delta t$, given by Eq. (9). We refer to the remainder, $\Omega_{t}$, as the “informatic heat” (IH). The propagator in the denominator, $\Pr(\mathbf{x}_{k},\lambda_{k}|\mathbf{x}_{k+1},\lambda_{k+1})$, is the exact same propagator of Eq. (2), where the jump process is reverted [i.e., the process starts at $\mathbf{x}_{k+1}$ at time $t=k\Delta t$ and ends up at $\mathbf{x}_{k}$ at time $(k+1)\Delta t$ under the protocol value $\lambda_{k+1}$].
@@
Therefore, if $\mathbf{D}$ and $\boldsymbol{\mu}$ are not related by a scalar, the informatic $\dot{\Sigma}_{t}$ is no longer related to the measurable $\dot{R}_{t}$, even on average. Consequently, the non-negativity of $\langle\dot{R}_{t}\rangle$ (i.e., the second law) can no longer be established based on trajectory statistics. Below we propose a definition of a nonequilibrium temperature that mends this discrepancy.
@@
We now impose the following physical conditions on Eq. (14): (i) To satisfy the second law, $\langle\dot{R}_{t}\rangle$ should be a non-negative functional of $\mathbf{J}$. (ii) To agree with Onsager’s theory [63, 64, 65], the right-hand side should reduce to a quadratic form of $\mathbf{J}$ near equilibrium and turn the second-law inequality into an equality for reversible processes ($\mathbf{J}\to\mathbf{0}$). (iii) The variable $\mathcal{T}$ should be a configuration-independent scalar. The simplest choice for the dissipation functional that satisfies these three conditions is $\langle\dot{R}_{t}\rangle=\int{\text{d}}\mathbf{x}(\mathbf{J}\cdot\mathbf{D}^{-1}\cdot\mathbf{J})/p$ [66]. To complete our construction, we solve the equation $\langle\dot{R}_{t}\rangle=\langle\dot{S}_{t}\rangle-\langle\dot{Q}_{t}\rangle/(k_{\mathrm{B}}\mathcal{T})$ for $k_{\mathrm{B}}\mathcal{T}$, and find the following drive-dependent generalized temperature:
@@
We have thus recovered the second law of thermodynamics for general Markovian overdamped Langevin dynamics (Eq. (2)).
Equation (15) is our central result. If ER holds, such that $\mathbf{D}^{-1}\cdot\boldsymbol{\mu}=(k_{\mathrm{B}}T)^{-1}$, we readily get $\mathcal{T}=T$. Otherwise, the generalized temperature is an intricate quantity. Computing it directly from Eq. (15) is a challenge. Below we propose how to estimate it from experiments (see, e.g., Eq. (18)). This is expected for nonequilibrium systems with arbitrarily complex $\mathbf{J}$, $\mathbf{D}$, $\boldsymbol{\mu}$, and $\mathbf{F}$. Furthermore, $\mathcal{T}$ depends on time via the protocol $\Lambda_{t}$ affecting all these quantities. Even for a constant protocol $\Lambda_{t}=\mathrm{const}$, $\mathcal{T}$ will change in time due to the instantaneous distribution $p(\mathbf{x},t)$ entering $\mathbf{J}$. Indeed, nonequilibrium steady states generally have $\mathbf{J}\neq\mathbf{0}$, as nonconservative forces and drift create persistent fluxes across scales. Despite these caveats, Eq. (15) should play an experimentally meaningful role as a generalized temperature in nonequilibrium thermodynamics [23, 16, 17, 18], as discussed below.
@@
which, combined with Eq. (9) and the relation $\dot{\Sigma}_{t}=\dot{S}_{t}+\dot{\Omega}_{t}$, gives the IEP rate as (cf. Ref. [67])
@@
Note that the correspondence holds for the means only and not for the full statistics, $\dot{Q}_{t}\neq-k_{\mathrm{B}}\mathcal{T}\dot{\Omega}_{t}$ and $\dot{R}_{t}\neq\dot{\Sigma}_{t}$. For example, the variances of the former are different, $\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}\cdot(\boldsymbol{\mu}\cdot\mathbf{F})p{\text{d}}\mathbf{x}$, whereas $\langle({\text{d}}Q_{t})^{2}\rangle_{\mathrm{c}}=2{\text{d}}t\int\mathbf{F}\cdot\mathbf{D}\cdot\mathbf{F}p{\text{d}}\mathbf{x}\neq(k_{\mathrm{B}}\mathcal{T})^{2}\langle({\text{d}}\Omega_{t})^{2}\rangle_{\mathrm{c}}$. (The subscript ${\mathrm{c}}$ denotes a cumulant.) Since the Jarzynski equality [38, 39] and Crooks fluctuation theorem [40] require statistics beyond the mean, they are not restored by the generalized $\mathcal{T}$. Thus, without the ER, the various FRs [37] remain related only to informatic quantities (the IEP and IH) [45] and not to thermodynamics (dissipation and heat).
@@
Nevertheless, generalized versions of the FDT and the Harada-Sasa relation can be derived in the absence of the ER. Here we outline these results; see the Supplemental Material for details [70]. Instead of considering the response to an impulsive linear potential gradient as was done in Ref. [68], we consider a perturbation of the force of the form $\delta\mathbf{F}(\mathbf{x},s)=-\delta h_{\mu}(s)\boldsymbol{\mu}^{-1}(\mathbf{x},s)\cdot\mathbf{D}(\mathbf{x},s)\cdot{\boldsymbol{\nabla}}x_{\mu}$ at time $s<t$, in response to the external perturbation $\delta\mathbf{h}(s)$. The difference between the response of the average force to such perturbation and its correlation with the position are found to be related to the thermodynamic heat [70],
@@
$\langle\dot{\Omega}_{s}\rangle=\lim_{t\to s^{+}}\left(\frac{{\text{d}}}{{\text{d}}s}\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}](\mathbf{X}_{t},t)\cdot\mathbf{X}_{s}\rangle\right.\\ \left.+\frac{\delta}{\delta\mathbf{h}(s)}\cdot\langle[(\boldsymbol{\mu}\cdot\mathbf{F})\cdot\mathbf{D}^{-1}](\mathbf{X}_{t},t)\rangle\right)\,.$ (21)